### v1

# Adaptive Canonicalization With Application to Invariant Anisotropic Geometric Networks

Abstract
Canonicalization is a widely used strategy in equivariant machine learning, enforcing symmetry in neural networks by mapping each input to a standard form. Yet, it often introduces discontinuities that can affect stability during training, limit generalization, and complicate universal approximation theorems. In this paper, we address this by introducing adaptive canonicalization, a general framework in which the canonicalization depends both on the input and the network. Specifically, we present the adaptive canonicalization based on prior maximization, where the standard form of the input is chosen to maximize the predictive confidence of the network. We prove that this construction yields continuous and symmetry-respecting models that admit universal approximation properties.
We propose two applications of our setting: (i) resolving eigenbasis ambiguities in spectral graph neural networks, and (ii) handling rotational symmetries in point clouds. We empirically validate our methods on molecular and protein classification, as well as point cloud classification tasks. Our adaptive canonicalization outperforms the three other common solutions to equivariant machine learning: data augmentation, standard canonicalization, and equivariant architectures.
1 Introduction
Equivariant machine learning [Gerken et al., 2023, Villar et al., 2021, Han et al., 2022, Keriven and Peyré, 2019] has been accentuated in geometric representation learning [Bronstein et al., 2017], motivated by the need to build models that respect symmetry inherent in data. For example, permutation equivariance in graphs [Gilmer et al., 2017, Zaheer et al., 2017, Xu et al., 2018], translation equivariance in images [LeCun and Bengio, 1998, Cohen and Welling, 2016a], and SO(3) or SE(3) equivariance for 3D objects and molecules [Thomas et al., 2018, Fuchs et al., 2020, Batzner et al., 2022, Satorras et al., 2021]. The symmetry is built into the method so that transforming the input induces a predictable transformation of the output. This inductive bias reduces sample complexity, curbs overfitting to arbitrary poses, and often improves robustness on distribution shifts where the same object appears in a different orientation or ordering [Kondor and Trivedi, 2018, Wang et al., 2022, Park et al., 2022, Bronstein et al., 2021, Bietti and Mairal, 2019, Kaba and Ravanbakhsh, 2023].
There are three principal approaches to handling symmetry in machine learning. The first involves designing equivariant architectures [Cohen and Welling, 2016b, Weiler et al., 2018a, Weiler and Cesa, 2019, Geiger and Smidt, 2022, Maron et al., 2019, Lippmann et al., 2024]: neural network layers are constructed to commute with the symmetry. The second approach is data augmentation, where each datapoint is presented to the model at an arbitrary pose [Chen et al., 2020, Brandstetter et al., 2022]. The third strategy is canonicalization [Kaba et al., 2023, Ma et al., 2023, 2024, Lim et al., 2022, 2023, Mondal et al., 2023, Lawrence et al., 2025, Sareen et al., 2025, Luo et al., 2022]: each input is mapped to a standard form and then processed by a non-equivariant network. Another less common approach to equivariant machine learning is frame averaging [Puny et al., 2021], which averages the network’s output over a set of input transformations.
A well-known problem in canonicalization is that in many cases it unavoidably leads to an end-to-end architecture which is discontinuous with respect to the input [Dym et al., 2024, Zhang et al., 2019a, Lim et al., 2022]. This inevitably leads to problems in stability during training and in generalization, as very similar inputs can lead to very different outputs [Dym et al., 2024, Tahmasebi and Jegelka, 2025a, b]. Moreover, the discontinuity of the network makes universal approximation properties less natural, as one approximates continuous symmetry preserving functions with discontinuous networks [Dym et al., 2024, Kaba et al., 2023, Wagstaff et al., 2022]. See App. A for extended related work.
Our Contribution.
In this paper, we show that the continuity problem in canonicalization can be solved if, instead of canonicalizing only as a function of the input, one defines a canonicalization that depends both on the input and the network. We propose such a general setting, which we call adaptive canonicalization, and show that it leads to continuous end-to-end models that respect the symmetries of the data and have universal approximation properties. Our theory does not only lead to superior theoretical properties w.r.t. standard canonicalization, but often also to superior empirical performance, specifically, in molecular, protein, and point cloud classification.
We focus on a specific class of adaptive canonicalizations that we call prior maximizers. To explain these methods, we offer the following illustrative example. Suppose that we would like to train a classifier of images into cats, dogs and horses. Suppose as well that each image $x$ can appear in the dataset in any orientation, i.e., as $\pi(\alpha)x$ for any $\alpha\in[0,2\pi]$ where $\pi(\alpha)$ is rotation by $\alpha$. One standard approach for respecting this symmetry is to design an equivariant architecture $\Theta$, which gives the same class probabilities to all rotations of the same image, i.e., $\Theta(\pi(\alpha)x)=\Theta(\pi(\alpha^{\prime})x)$ for any to angles $\alpha,\alpha^{\prime}$. Another simple approach for improving the classifier is to train a symmetryless network $\Psi$, and augment the training set with random rotations $\pi(\alpha)x$ for each input $x$. Yet another standard approach is to canonicalize the input, namely, to rotate each input image $x$ by an angle $\beta_{x}$ that depends on $x$ in such a way that all rotated versions of the same image would have the exact same standard form, i.e., $\pi(\beta_{\pi(\alpha)x})\pi(\alpha)x=\pi(\beta_{\pi(\alpha^{\prime})x})\pi(\alpha^{\prime})x$ for any two angles $\alpha,\alpha^{\prime}$. Then, the canonicalized image $\pi(\beta_{x})x$ is plugged into a standard symmetryless neural network $\Psi$, and the end-to-end architecture $\Psi(\pi(\beta_{x})x)$ is guaranteed to be invariant to rotations. We propose a fourth approach, where the canonicalized rotation $\pi(\beta_{x,\Psi})$ depends both on the image $x$ and on the (symmetryless) neural network $\Psi$.
Figure 1: Illustration of prior maximization adaptive canonicalization in classification. The adaptive canonicalization optimizes the transformations $\beta_{x,\Psi_{j}}$ of the inputs $x$ to the classifiers $\Psi_{j}$, while, during training, $\Psi_{j}$ are simultaneously trained w.r.t. the adaptively canonicalized inputs $\pi(\beta_{x,\Psi_{j}})x$.
To motivate this approach, consider a neural network $\Psi$ which, by virtue of being symmetriless, may perform better on some orientations of $x$ than others. For illustration, it is easier for humans to detect an image as a horse if its limbs point downwards. Suppose that $\Psi(x)=(\Psi_{j}(x))_{j=1}^{3}=(\Psi_{dog}(x),\Psi_{cat}(x),\Psi_{horse}(x))$ is a sequence of binary classifiers with values in $[0,1]$ each. The output of $\Psi$ is defined to be the class with highest probability. Suppose moreover that for each $j$, the network $\Psi_{j}$ is granted the ability to rotated $x$ freely, and probe the output $\Psi_{j}(\pi(\alpha)x)$ for each $\alpha$. The network then chooses the orientation $\alpha_{*}$ such that $\Psi_{j}(\pi(\alpha_{*})x)$ is maximized. As an analogy, one can imagine an image on a piece of paper being handed at a random orientation to a person with visual system $\Psi$. To detect if there is a horse in the image, the person would rotate the paper, searching for an orientation which looks like a horse. Namely, if there is an orientation $\alpha_{*}$ where $\Psi_{horse}(\pi(\alpha_{*})x)$ is high then there is a horse in the image, and otherwise there is non. This process would be repeated for all other classes, and eventually the image would be classified as the $j_{*}$ such that $\max_{\alpha}\Psi_{j_{*}}(\pi(\alpha)x)$ is greater than $\max_{\alpha}\Psi_{j}(\pi(\alpha)x)$ for all other $j\neq j_{*}$. This is the process that we call prior maximization adaptive canonicalization.
We note that the inputs are adaptively canonicalized also during training, so $\Psi$ needs not learn to respect any symmetry on its own. In fact, $\Psi$ can benefit from being symmetryless. For example, it may search for “horse head” patterns only diagonally above where it detects “horse limbs” patterns, and rely on the prior maximization to orient horses accordingly.
We show that adaptive canonicalization leads to a continuous symmetry preserving end-to-end classifier that can approximate any symmetry preserving continuous function when $\Psi$ are non-equivariant neural networks. As an application, we propose adaptive canonicalization methods for 1) spectral graph neural networks, where the symmetry is in the choice of the eigenbasis of the graph shift operator, and 2) point clouds, with rotation symmetries. We show that adaptive canonicalization in these cases outperforms both standard canonicalization and equivariant networks, as well as augmentation methods. See Fig. 1 for an illustration of prior maximization.
2 Adaptive Canonicalization
In this section, we develop the general theory of adaptive canonicalization, and prove that it leads to continuous symmetry preserving networks with universal approximation properties.
2.1 Basic Definitions and Background
The function that maps each input $x$ to the output $f(x)$ is denoted by $x\mapsto f(x)$. The free variable of a univariate function is denoted by $(\cdot)$, and by $(\cdot,\cdot\cdot)$ for a function of two variables. For example, the function $(x,y)\mapsto\sin(x)\exp(y)$ is also denoted by $\sin(\cdot)\exp(\cdot\cdot)$. We denote the infinity norm of $x=(x_{d})_{d=1}^{D}\in\mathbb{R}^{D}$ by $\left\lvert x\right\rvert:=\max_{1\leq d\leq D}\left\lvert x_{d}\right\rvert$. We define the infinity norm of a continuous function $f:\mathcal{K}\rightarrow\mathbb{R}^{D}$ over a topological space $\mathcal{K}$ by $\|f\|_{\infty}=\sup_{x\in\mathcal{K}}\left\lvert f(x)\right\rvert$. If $\|f-y\|_{\infty}<\epsilon$ we say that $y$ approximates $f$ uniformly up to error $\epsilon$. The set of all subsets of a set $\mathcal{K}$, i.e., the power set, is denoted by $2^{\mathcal{K}}$. When defining general metric spaces, we allow the distance between points to be $\infty$. This does not affect most of the common properties of metric spaces (see Burago et al. [2001]).
Function Spaces.
We develop the definitions of adaptive canonicalization in general locally compact Hausdorff spaces. Two important examples of such a space is a compact metric space or $\mathbb{R}^{J}$.
Definition 1.
Let $\mathcal{K}$ be a locally compact Hausdorff space, and $D\in\mathbb{N}$.
  • A function $f:\mathcal{K}\rightarrow\mathbb{R}^{D}$ is said to vanish at infinity is for every $\epsilon>0$ there exists a compact set $K\subset\mathcal{K}$ such that $\left\lvert f(x)\right\rvert<\epsilon$ for every $x\in\mathcal{K}\setminus K$.
  • The space of all continuous functions $f:\mathcal{K}\rightarrow\mathbb{R}^{D}$ that vanish at infinity, with the supremum norm $\|f\|_{\infty}=\max_{x\in\mathcal{K}}\left\lvert f(x)\right\rvert$ is denoted by $C_{0}(\mathcal{K},\mathbb{R}^{D})$.
In adaptive canonicalization, we consider families of continuous functions where the $\epsilon-\delta$ formulation of continuity is uniform over the whole family, as defined next.
Definition 2.
Let $\mathcal{X}$ and $\mathcal{Y}$ be two metric spaces with metrics $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$ respectively. A family $\mathcal{F}$ of function $f:\mathcal{X}\rightarrow\mathcal{Y}$ is called equicontinuous if for every $x\in\mathcal{X}$ and every $\epsilon>0$, there exists $\delta>0$ such that every $z\in\mathcal{X}$ which satisfies $d_{\mathcal{X}}(x,z)<\delta$ also satisfies
$\forall f\in\mathcal{F}:\quad d_{\mathcal{Y}}(f(x),f(z))<\epsilon.$
Universal Approximation.
Universal approximation theorems (UAT) state that any continuous function over some topological space can be approximated by a neural network. In such a case, the neural networks are said to be universal approximators, as defined next.
Definition 3.
Let $\mathcal{K}$ be a locally compact Hausdorff space and $D\in\mathbb{N}$. A set of continuous functions $\mathcal{N}(\mathcal{K},\mathbb{R}^{D})\subset C_{0}(\mathcal{K},\mathbb{R}^{D})$ is said to be a universal approximator of $C_{0}(\mathcal{K},\mathbb{R}^{D})$ if for every $f\in C_{0}(\mathcal{K},\mathbb{R}^{D})$ and $\epsilon>0$ there is a function $\theta\in\mathcal{N}(\mathcal{K},\mathbb{R}^{D})$ such that
$\forall x\in\mathcal{K}:\quad\left\lvert f(x)-\theta(x)\right\rvert<\epsilon.$
In the above definition, we interpret $\mathcal{N}(\mathcal{K},\mathbb{R}^{D})$ as a space of neural networks. A UAT is hence any theorem which shows that some set of neural networks is a universal approximator. Two examples of UATs are: 1) multilayer perceptrons (MLP) are universal approximators of $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ for compact subset $\mathcal{K}$ of the Euclidean space ${\mathbb{R}}^{D}$ [Hornik et al., 1989, Cybenko, 1989], and 2) DeepSets [Zaheer et al., 2017] are universal approximators of continuous functions from multi-sets to ${\mathbb{R}}^{d}$. See App. B for more details.
2.2 Adaptive Canonicalization
In the general setting of adaptive canonicalization, we have a domain of inputs $\mathcal{G}$ which need not have any structure apart for being a set, e.g., the set of graphs. We consider continuous functions $f:\mathcal{K}\rightarrow\mathbb{R}^{L}$ over a “nice” domain $\mathcal{K}$, e.g., $\mathcal{K}={\mathbb{R}}^{J}$. Such functions can be approximated by neural networks. We then pull-back $f$ to be a function from $\mathcal{G}$ to $\mathbb{R}^{L}$ using a mapping $\rho_{f}:\mathcal{G}\mapsto\mathcal{K}$ that depends on (is adapted to) $f$. Namely, we consider $f(\rho_{f}(\cdot)):\mathcal{G}\rightarrow\mathbb{R}^{D}$. The following definitions assure that such a setting leads to functions with nice properties, as we show in subsequent sections.
Definition 4.
Let $\mathcal{K}$ be a locally compact Hausdorff space, $\mathcal{G}$ be a set, and $D\in\mathbb{N}$. A mapping $\rho=\rho_{(\cdot)}(\cdot\cdot):C_{0}(\mathcal{K},\mathbb{R}^{D})\times\mathcal{G}\rightarrow\mathcal{K}$, $(f,g)\mapsto\rho_{f}(g)$, is called an adaptive canonicalization if the set of functions
$\{f\mapsto f\circ\rho_{f}(g)\ |\ g\in\mathcal{G}\}$
is equicontinuous (as functions $C_{0}(\mathcal{K},\mathbb{R})\rightarrow\mathbb{R}^{D}$). Here, $f\circ\rho_{f}(g):=f(\rho_{f}(g))$.
Next, we define the function space that we would like to approximate using adaptive canonicalization.
Definition 5.
Let $\mathcal{K}$ be a locally compact Hausdorff space, $\mathcal{G}$ be a set, and $D\in\mathbb{N}$. Let $\rho$ be an adaptive canonicalization, and let $f\in C_{0}(\mathcal{K},\mathbb{R}^{D})$. The function
$f\circ\rho_{f}:\mathcal{G}\rightarrow{\mathbb{R}}^{D},\quad g\mapsto f\big(\rho_{f}(g)\big)$
is called an adaptive canonicalized continuous function, or a canonicalized function in short.
In Sec. 2.4, we show that for an important class of adaptive canonicalizations the set of adaptive canonicalized continuous functions is exactly the set of all symmetry preserving continuous functions.
It is now direct to prove the following universal approximation theorem.
Theorem 6 (Universal approximation of adaptive canonicalized functions).
Let $\mathcal{N}(\mathcal{K},\mathbb{R}^{D})$ be a universal approximator of $C_{0}(\mathcal{K},\mathbb{R}^{D})$, and $f\circ\rho_{f}$ an adaptive canonicalized continuous function. Then, for every $\epsilon>0$, there exists a network $\theta\in\mathcal{N}(\mathcal{K},\mathbb{R}^{D})$ such that for every $g\in\mathcal{G}$
$\left\lvert f\circ\rho_{f}(g)-\theta\circ\rho_{\theta}(g)\right\rvert<\epsilon.$
Proof.
Let $\epsilon>0$. By Def. 4, there exists $\delta>0$ such that
$\forall y\in C_{0}(\mathcal{K},\mathbb{R}^{D}):\ \|f-y\|_{\infty}<\delta\Rightarrow\Big(\forall g\in\mathcal{G}:\ \left\lvert f\circ\rho_{f}(g)-y\circ\rho_{y}(g)\right\rvert<\epsilon\Big).$ (1)
By the the universal approximation property, there exists a network $\theta$ such that $\|f-\theta\|_{\infty}<\delta$. Hence, by (1), $\quad\forall g\in\mathcal{G}:\ \left\lvert f\circ\rho_{f}(g)-\theta\circ\rho_{\theta}(g)\right\rvert<\epsilon$. ∎
2.3 Prior Maximization Adaptive Canonicalization
Prior maximization is a special case of adaptive canonicalization, where $\rho_{f}$ is chosen to maximize some prior on the output of $f$. The maximization is done over a space of transformations $\kappa_{u}:\mathcal{G}\rightarrow\mathcal{K}$ parameteried by $u$, i.e., maximizing the prior of $f(\kappa_{u}(g))$ w.r.t. $u$.
Let $\mathcal{U}$ be metric space, and for every $u\in\mathcal{U}$, let
$\kappa_{(\cdot)}(\cdot\cdot):\mathcal{U}\times\mathcal{G}\rightarrow\mathcal{K},\quad(u,g)\mapsto\kappa_{u}(g)\in\mathcal{K}.$
Suppose that $\kappa_{u}(g)$ is continuous in $u$ for every $g\in\mathcal{G}$. We call $\kappa$ a transformation family, transforming objects in $\mathcal{G}$ into points in $\mathcal{K}$, where different $u\in\mathcal{U}$ define different transformations. Let $H=(h_{1},\ldots,h_{D})$, where $h_{d}:\mathbb{R}\rightarrow\mathbb{R}$ for each $d$, be a sequence of continuous monotonic functions, that we call the ensemble of priors. We call each $h_{d}$ a prior. We denote $H\circ f:=(h_{d}\circ f_{d})_{d=1}^{D}$.
For every $f=(f_{1},\ldots,f_{D})\in C_{0}(\mathcal{K},\mathbb{R}^{D})$, $g\in\mathcal{G}$ and $d$, assume that $h_{d}\circ f_{d}(\kappa_{(\cdot)}(g))$ attains a maximum in $\mathcal{U}$. This is the case for example when $\mathcal{U}$ is compact. Define
$\rho_{f}(g)=\big(\rho^{d}_{f_{d}}(g)\big)_{d=1}^{D}:=\Big(\big\{\kappa_{u_{*}}(g)\ \big|\ h_{d}\circ f_{d}\big(\kappa_{u_{*}}(g)\big)=\max_{u\in\mathcal{U}}h_{d}\circ f_{d}\big(\kappa_{u}(g)\big)\big\}\Big)_{d=1}^{D}\in\big(2^{\mathcal{K}}\big)^{D}.$ (2)
Note that $\rho_{f}:\mathcal{G}\rightarrow(2^{\mathcal{K}})^{D}$. By abuse of notation, we also denote by $\rho_{f}$ the mapping that returns some arbitrary sequence of points $(x_{d}\in\rho^{d}_{f_{d}}(g))_{d=1}^{D}\in\mathcal{K}^{D}$ for each $g\in\mathcal{G}$. The choice of the specific point in $\rho^{d}_{f_{d}}(g)$ does not affect the analysis. We interpret $\rho_{f}$ as a function that takes an input $g$ and canonicalize it separately with respect to each output channel $f_{d}$, adaptively to $f_{d}$.
When used for classification, we interpret each output channel $f_{d}\circ\rho_{f_{d}}^{d}(g)\in[0,1]$ as a binary classifier, i.e., representing the probability of $g$ being in class $d$ vs. not being in class $d$. This multiclass classification setting is called one vs. rest, where a standard loss is a sum of $D$ binary cross-entropies [Rifkin and Klautau, 2004, Galar et al., 2011, Allwein et al., 2000].
Definition 7.
Consider the above setting. The mapping $\rho$ defined by (2) is called prior maximization. If in addition $\mathcal{G}$ has a metric such that for every $f\in C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ the family $\{g\mapsto f(\kappa_{u}(g))\}_{u\in\mathcal{U}}$ is equicontinuous, $\rho$ is called continuous prior maximization.
In Thm. 8 we show that prior maximization is indeed adaptive canonicalization.
Note that the condition of $g\mapsto f(\kappa_{u}(g))$ being equicontinuous is satisfied for well known settings of equivariant machine learning. For example, let $\mathcal{U}=\mathcal{SO}(3)$ be the space of 3D rotations, and $\mathcal{G}=\mathcal{K}=\mathcal{B}^{N}$ the set of sequences of $N$ points in the 3D unit ball $\mathcal{B}$, i.e., the space of point clouds. We consider the rotation $g\mapsto\kappa_{u}(g)$ of the point cloud $g$ by $u\in\mathcal{U}$. Since $\mathcal{G}$ and $\mathcal{U}$ are compact metric spaces, and $(g,u)\mapsto\kappa_{u}(g)$ is continuous, $\kappa$ and $f$ must be uniformly continuous. Hence, $\{g\mapsto f(\kappa_{u}(g))\}_{u\in\mathcal{U}}$ is equicontinuous. In fact, whenever $\mathcal{G}$ is compact and $\kappa$ continuous w.r.t. $(u,g)$, it is automatically also uniformly continuous, so $\rho$ is a continuous prior maximization. See App. C for additional examples of continuous prior maximization.
Properties of Prior Maximization.
Theorem 8.
In prior maximization, each $\rho^{d}:C_{0}(\mathcal{K},{\mathbb{R}})\times\mathcal{G}\rightarrow\mathcal{K}$ is adaptive canonicalization.
Proof.
Consider without loss of generality the case where the output dimension is $D=1$. Since the specific values of $H=h_{1}$ do not matter, only if it is ascending or descending, without loss of generality suppose $H(x)=x$, in which case prior maximization maximizes directly the output of $f\circ\kappa_{u}(g)$ with respect to $u$. Consider an arbitrary maximizer $\rho_{f}(g)\in{\rm arg}\max_{u\in\mathcal{U}}f(\kappa_{u}(g))$ for each $f\in C_{0}(\mathcal{K},{\mathbb{R}})$. The choice of the maximizer does not affect the analysis. Now, if $f,y\in C_{0}(\mathcal{K},{\mathbb{R}})$ satisfy $\|f-y\|_{\infty}<\epsilon$, then also for every $u\in\mathcal{U}$, $\left\lvert f(\kappa_{u}(g))-y(\kappa_{u}(g))\right\rvert<\epsilon$. Let $u_{0}\in\mathcal{U}$ be a maximizer of $f(\kappa_{u}(g))$. We have $y(\kappa_{u_{0}}(g))>f(\kappa_{u_{0}}(g))-\epsilon$, so
$\max_{u}y(\kappa_{u}(g))>\max_{u}f(\kappa_{u}(g))-\epsilon.$
Similarly, we have $\max_{u}f(\kappa_{u}(g))>\max_{u}y(\kappa_{u}(g))-\epsilon$. Together,
$\left\lvert\max_{u\in\mathcal{U}}f\circ\kappa_{u}(g)-\max_{u\in\mathcal{U}}y\circ\kappa_{u}(g)\right\rvert<\epsilon.$ (3)
Hence, $f\mapsto f\circ\rho_{f}(g)$ is Lipschitz continuous with Lipschitz constant 1 for every $g\in\mathcal{G}$, and therefore equicontinuous over the parameter $g\in\mathcal{G}$. ∎
This immediately gives a universal approximation theorem for prior maximization as a corollary of Thm. 6. Moreover, we can show that continuous prior maximization gives functions continuous in $\mathcal{G}$. This is one of the main distinctions between prior maximization and standard canonicalization.
Theorem 9.
Consider a continuous prior maximization $\rho$ (Def. 7). Then, $f\circ\rho_{f}$ is continuous.
Proof.
Let $\epsilon>0$. For every $g\in\mathcal{G}$ there is $\delta=\delta_{\epsilon,g}>0$ such that for every $g^{\prime}\in\mathcal{G}$ with $d(g,g^{\prime})<\delta$ and every $u\in\mathcal{U}$ we have $\left\lvert f(\kappa_{u}(g))-f(\kappa_{u}(g^{\prime}))\right\rvert<\epsilon$. Now, by the same argument as in (3),
$\left\lvert\max_{u}f(\kappa_{u}(g))-\max_{u}f(\kappa_{u}(g^{\prime}))\right\rvert<\epsilon.$
2.4 Symmetry Preserving Prior Maximization
[…]
Definition 10.
Consider the above setting and a continuous prior maximization $\rho$. Let $P:\mathcal{G}\rightarrow\mathcal{K}$ be continuous, and suppose that the transformation family $\kappa$ is of the form $\kappa_{u}=P\circ\pi(u)$. We call $\kappa$ a symmetry preserving transformation family, and $\rho$ a symmetry preserving prior maximization.
[…]
Definition 11.
We call a function $Q:\mathcal{G}\rightarrow{\mathbb{R}}^{D}$ continuous symmetry preserving if there exists $F:\mathcal{K}\rightarrow{\mathbb{R}}^{D}$ in $C_{0}(\mathcal{K};{\mathbb{R}}^{D})$ such that for all $u\in\mathcal{U}$ and $g\in\mathcal{G}$, $\quad Q(g)=F(P(\pi(u)(g)))$.
When $\mathcal{K}=\mathcal{G}=\mathcal{G}_{1}$ and $P$ is the identity, a symmetry preserving continuous function is a continuous function which is invariant to the action of $u$, i.e., the classical case in equivariant machine learning.
Properties of Symmetry Preserving Prior Maximization.
We already know by Thm. 9 that $f\circ\rho_{f}$ is a continuous function when $\rho$ is a symmetry preserving prior maximization. We next show that the set of functions of the form $f\circ\rho_{f}$ exhaust the space of all continuous symmetry preserving functions.
Theorem 12.
Let $\rho$ be a symmetry preserving adaptive canonicalization. Then,
  1. 1.
    Any continuous symmetry preserving function can be written as $f\circ\rho_{f}$ for some $f\in C_{0}(\mathcal{K},{\mathbb{R}}^{D})$.
  2. 2.
    For any $f\in C_{0}(\mathcal{K};{\mathbb{R}}^{D})$, the function $f\circ\rho_{f}:\mathcal{G}\rightarrow{\mathbb{R}}^{D}$ is continuous symmetry preserving.
Proof.
Without loss of generality, consider the case $D=1$ and $H(x)=x$.
  1. 1.
    For $Q(g)=F(P(\pi(u)(g)))$, take $f=F$. Then, by definition of symmetry preservation, for every $u\in\mathcal{U}$: $\quad f\circ\rho_{f}(g)=\max_{v}F(P\circ\pi(v)g)=F(P\circ\pi(u)g)=Q(g)$.
  2. 2.
    For any $u=u_{i}\in\mathcal{U}_{i}$, since $\pi_{i}(u_{i})$ is an action, for any $g=g_{j}\in\mathcal{G}_{j}$,
    $f\circ\rho_{f}(\pi(u)g)=\max_{v_{j}\in\mathcal{U}_{j}}f(P\circ\pi_{j}(v_{j})\pi_{i}(u_{i})g_{j})=\max_{v_{j}\in\mathcal{U}_{j}}f(P\circ\pi_{j}(v_{j})g_{j})=f\circ\rho_{f}(g).$
This leads to the following UAT.
Theorem 13.
Consider a symmetry preserving prior maximization and let $\mathcal{N}(\mathcal{K},{\mathbb{R}}^{D})$ be a universal approximator of $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ . Then, any continuous symmetry preserving function can be approximated uniformly by $\theta\circ\rho_{\theta}$ for some network $\theta\in\mathcal{N}(\mathcal{K},{\mathbb{R}}^{D})$.
3 Application of Adaptive Canonicalization to Anisotropic Geometric Networks
In this section, we propose two architectures based on adaptive canonicalization which can be interpreted as anisotropic. First, a spectral graph neural network (GNN) which is sensitive to directionality within eigenspaces. Then, a 3D point cloud network which is sensitive to 3D directions.
Basic Notations for Graphs and Vectors.
We denote by $\mathbb{N}_{0}$ the set of nonnegative integers. We denote matrices and 2D arrays by boldface capital letters, e.g., ${\mathbf{B}}\in{\mathbb{R}}^{N\times T}$. We denote $N]=\{1,\ldots,N\}$ for $N\in\mathbb{N}$. We denote by ${\mathbf{B}}(:,j)$ and ${\mathbf{B}}(j,:)$ the $j$’th column and row of the matrix $\mathbf{B}$} respectively. A graph is denoted by $G=([N],{\mathbf{A}},{\mathbf{S}})$, where $[N]$ is the set of $N$ vertices, ${\mathbf{A}}\in{\mathbb{R}}^{N\times N}$ is the adjacency matrix, and ${\mathbf{S}}\in{\mathbb{R}}^{N\times T}$ is an array representing the signal. Namely, ${\mathbf{S}}(n,:)\in{\mathbb{R}}^{T}$ is the feature at node $n$. A graph shift operator (GSO) is a self-adjoint operator that respects in some sense the graph’s connectivity, e.g., a graph Laplacian or the adjacency matrix.
3.1 Anisotropic nonlinear spectral filters
[…]
In the above setting, the galaxies $\mathcal{G}_{j}$ are defined as follows. Each galaxy is indexed by $j=(N,{\mathbf{M}}):=(N;M_{1},\ldots,M_{B})$ where $N\in\mathbb{N}$ is the size of the graph, and $M_{k}\in\mathbb{N}_{0}$ for $k\in[B]$ are the dimensions of the bands, satisfying $\sum_{k}M_{k}\leq N$. Each $\mathcal{G}_{N,{\mathbf{M}}}$ is the space of pairs of a signal $\mathbf{S}\in{\mathbb{R}}^{N\times T}$ and a sequence of $K$ orthogonal matrices of height $N$ and widths $M_{1},\ldots,M_{B}$. We denote the matrix with dimension 0 by $\emptyset$. For two signals $\mathbf{S},\mathbf{Q}$ and orthogonal bases sequences $\mathbf{X},{\mathbf{Y}}$, the metric $d\big(({\mathbf{X}},\mathbf{S}),({\mathbf{Y}},\mathbf{Q})\big)$ is defined to satisfy
$d({\mathbf{X}},{\mathbf{Y}})^{2}=\|\mathbf{S}-\mathbf{Q}\|_{2}^{2}+\sum_{k=1}^{B}\sum_{j=1}^{M_{k}}\|{\mathbf{X}}_{k}[:,j]-{\mathbf{Y}}_{k}[:,j]\|_{2}^{2}=\|\mathbf{S}-\mathbf{Q}\|_{2}^{2}+\|\mathbf{X}-\mathbf{Y}\|_{2}^{2},$
where for $\mathbf{z}=(z_{n})_{n}$, $\|\mathbf{z}\|_{2}^{2}=\sum_{n}z_{n}^{2}$. By definition, the distance between any two points from two different galaxies is $\infty$.
The transfomation family $\kappa_{{\mathbf{U}}}$ apply unitary operators ${\mathbf{U}}_{k}\in\mathcal{O}(M_{k})$ form the right on each ${\mathbf{X}}(:,k)$, where $\mathcal{O}(M_{k})$ is the group of orthogonal matrices in ${\mathbb{R}}^{M_{k}\times M_{k}}$. Namely, $\mathcal{U}_{N;\mathbf{M}}:=\prod_{k=1}^{B}\mathcal{O}(M_{k})$. We now define the operator $C$ that computes the spectral coefficients of the signal $\mathbf{S}$ with respect to the spectral basis $\mathbf{X}$. Namely,
$C({\mathbf{X}},\mathbf{S})=\big(C_{k}({\mathbf{X}},\mathbf{S})\big)_{k=1}^{B}:=({\mathbf{X}}(:,k)^{\top}\mathbf{S}\in{\mathbb{R}}^{M_{k}\times T})_{k=1}^{B}.$
In anisotropic nonlinear spectral filters (A-NLSF), we consider a symmetryless neural network $\Psi$ that operates on the space of spectral coefficients of the signal, e.g., a multilayer perceptron (MLP). To define $\Psi$ consistently, over a Euclidean space of fixed dimension, we extend or truncate the sequence of spectral coefficients as follows. Let $J_{1},\ldots,J_{B}$ be a predefined sequence of integers, and denote $J=\sum_{k}J_{k}$. We define $P$ as the mapping that takes $({\mathbf{X}},\mathbf{S})$ as input, first compute the spectral coefficients $\big(C_{k}({\mathbf{X}},\mathbf{S})\big)_{k=1}^{B}$, and then truncates or pads with zeros each $C_{k}({\mathbf{X}},\mathbf{S})$ to be in ${\mathbb{R}}^{J_{k}\times T}$. Namely, $P({\mathbf{X}},\mathbf{S})=\big(P_{k}({\mathbf{X}},\mathbf{S})\in{\mathbb{R}}^{J_{k}\times T}\big)_{k=1}^{B}\in{\mathbb{R}}^{J\times T}$. Here, the matrix $\emptyset$ is padded to the zero matrix $\mathbf{0}\in{\mathbb{R}}^{J_{k}\times T}$. Hence, the symmetryless network $\Psi$ maps ${\mathbb{R}}^{J\times T}$ to ${\mathbb{R}}^{D}$. A more detailed derivation of the construction is given in App. E.1.
We call this architecture anisotropic for the following reason. Consider for example the grid graph with $N\times N$ vertices. Since spectral filters are based on functional calculus (see App. D), they are invariant to graph automorphism, and hence to rotations. This means that spectral filters treat the $x$ and $y$ axes equally, and any filter is isotropic in the spatial domain. On the other hand, our symmetryless network $\Psi$ can operate differently on the $x$ and $y$ axes, and we can implement general directional filters on images with $\Psi$. In the general case, $\Psi$ can operate differently on Fourier modes from the same eigenspace, which we interpret as directions within the eigenspace, while standard GNNs cannot. See App. G.2 for more details.
3.2 Anisotropic Point Cloud Networks
Here, we present a point cloud network which is a combination of an equivariant network with adaptive canonicalization. Namely, we consider a permutation invariant network $\Psi$ like DeepSet [Zaheer et al., 2017] or DGCNN [Wang et al., 2019], and to attain 3D rotation invariance in addition we incorporate prior maximization. Together, the method is invariant both to permutations and rotations. We call this method anisotropic since $\Psi$ does not respect the rotation symmetries, and is hence sensitive to directions in the $x,y,z$ space.
We restrict the analysis to multi-sets of a fixed number of points $N$. Multisets are sets where repetitions of elements are allowed. Here, we formally define a multi-set as an equivalence class of arrays up to permutation. To define this, let $\mathcal{S}_{N}$ be the symmetric group of $N$ elements, i.e., the group of permutations. given $s\in\mathcal{S}$ and ${\mathbf{X}}\in{\mathbb{R}}^{N\times J}$, let $\rho(s){\mathbf{X}}$ be the permutation that changes the order of the rows $\mathbf{X}$} according to $s$. We say that ${\mathbf{X}}\sim{\mathbf{Y}}$ if there is $s\in\mathcal{S}_{N}$ such that ${\mathbf{X}}=\rho(s){\mathbf{Y}}$. The equivalence class $\mathbf{X}$}] is defined as $\{{\mathbf{Y}}\in{\mathbb{R}}^{N\times J}\ |\ {\mathbf{Y}}\sim{\mathbf{X}}\}$, and the space of equivalence classes, also called the quotient space, is denoted by $({\mathbb{R}}^{N\times J}/\sim):=\{[{\mathbf{X}}]\ |\ {\mathbf{X}}\in{\mathbb{R}}^{N\times J}\}$. We identify the space $({\mathbb{R}}^{N\times J}/\sim)$ with the space of multisets.
In App. B.2 we show that the quotient space has a natural metric. We hence take $\mathcal{G}=\mathcal{K}$ consisting of a single galaxy $\mathcal{G}_{N}=({\mathbb{R}}^{N\times J}/\sim)$. We moreover show in App. B.2 that any universal approximator of permutation invariant functions in $C_{0}({\mathbb{R}}^{N\times J},{\mathbb{R}}^{D})$, e.g., DeepSet [Zaheer et al., 2017], canonically gives a universal approximator of general continuous functions in $C_{0}({\mathbb{R}}^{N\times J}/\sim,{\mathbb{R}}^{D})$.
The symmetry in our adaptive canonicalization is 3D rotations $\pi(u)\in{\mathbb{R}}^{3\times 3}$, where $u$ is in the rotation group $\mathcal{SO}(3)$. Namely, we consider $J=3$, and rotated the rows of ${\mathbf{X}}\in{\mathbb{R}}^{N\times 3}$ via ${\mathbf{X}}\pi(u^{-1})$. We take $P$ as the identity. Note that this construction can be easily extended to multisets of arbitrary sizes, by considering the galaxies $\mathcal{G}_{N}={\mathbb{R}}^{N\times 3}/\sim$ and groups $\mathcal{U}_{N}=\mathcal{S}_{N}$ for all $N\in\mathbb{N}$. For details on the theoretical construction see App. B.2, and for details on the architecture App. E.2.
Table 1: Classification performance on grid signal orientation task and graph classification benchmarks from TUDataset. The highest accuracy in   and the second highest in  .
Toy Example TUDataset
MUTAG PTC ENZYMES PROTEINS NCI1
[…]
4 Experiments
We evaluate the anisotropic nonlinear spectral filters (Sec. 3.1) on toy problems and graph classification, and test the anisotropic point cloud network (Sec. 3.2) on point cloud classification. The experimental details, including experimental setups and hyperparameters, are in App. F.
4.1 Experimental Evaluation of Anisotropic Nonlinear Spectral Filters
Illustrative Toy Problems: Grid Signal Orientation Task.
To showcase the effectiveness of A-NLSF, we study a toy classification task on a grid-split channel orientation. We consider a square grid on the torus, and each node has two channels. In addition, the grid is further partitioned vertically into two equal disjoint halves. Channel 1 is nonzero only on the left half, and Channel 2 is only on the right half. In class 0, both channels are 1-frequency along $x$, and in class 1, Channel 1 is 1-frequency along $x$ and Channel 2 is 1-frequency along $y$. The task is to decide if the frequency at the two channels is in the same orientation. See App. F for further details.
We compare A-NLSF with the following baselines: (i) MLP, (ii) GCN [Kipf, 2016], (iii) GAT [Veličković et al., 2017], (iv) GIN [Xu et al., 2018], (v) ChebNet [Defferrard et al., 2016a], (vi) NLSF [Lin et al., 2024a], and (vii) S2GNN [Geisler et al., 2024]. In addition, we test canonicalization baselines by combining GIN with frame averaging (FA) [Puny et al., 2021] and orthogonalized axis projection (OAP) [Ma et al., 2024]. Tab. LABEL:tab:medium_size_graph_classification reports the classification results. We see that competing methods remain at chance level while A-NLSF achieves high accuracy by adaptively resolving ambiguities, showing the advantage for orientation-sensitive learning on disjoint supports.
Graph Classification on TUDataset.
We further evaluate A-NLSF on graph classification benchmarks from TUDataset [Morris et al., 2020]: MUTAG, PTC, ENZYMES, PROTEINS, and NCI1, and follow the experimental setup [Ma et al., 2019, Ying et al., 2018, Zhang et al., 2019b] (see App. F). We compare with the same baselines as in the grid signal orientation tasks. Tab. LABEL:tab:medium_size_graph_classification summarizes the classification performance. Canonicalization baselines generally improve over GIN. Notably, we observe that A-NLSF outperforms competing baselines, suggesting that our AC provides more informative representations compared to a fixed, precomputed canonical form or isotropic filters.
Table 2: Molecular and protein classification performance on OGB datasets.
ogbg-molhiv ogbg-molpcba ogbg-ppa
AUROC $\uparrow$ Avg. Precision $\uparrow$ Accuracy $\uparrow$
[…]
Molecular Classification on OGB Datasets.
To further assess the effectiveness of A-NLSF, we evaluate on large-scale molecular and protein benchmarks from Open Graph Benchmark (OGB) [Hu et al., 2020]: ogbg-molhiv, ogbg-molpcba, and ogbg-ppa. We compare with GCN, GIN, GatedGCN [Bresson and Laurent, 2017], PNA [Corso et al., 2020], GraphTrans [Wu et al., 2021], SAT [Chen et al., 2022], GPS [Rampášek et al., 2022], SAN [Kreuzer et al., 2021], and the canonicalization method OAP+GatedGCN. The molecular classification results are reported in Tab. 2, showing that our method constructs robust molecular and protein representations, leading to improved generalization in classification.
4.2 Experimental Evaluation of Anisotropic Point Cloud Networks
Table 3: Classification results on ModelNet40.
Accuracy
PointNet 74.7
DGCNN 88.6
PointNet-Aug 76.4
DGCNN-Aug 89.3
VN-PointNet 77.2
[…]
CN-PointNet 79.7
CN-DGCNN 90.0
AC-PointNet 81.3
[…]
To evaluate adaptive canonicalization on point cloud classification, we test ModelNet40 [Wu et al., 2015]. The dataset consists of 12,311 shapes from 40 categories, with 9,843 samples for training and 2,468 for testing. We build on two point cloud architectures, PointNet [Qi et al., 2017] and DGCNN [Wang et al., 2019], and apply adaptive canonicalization into their pipeline, denoted respectively AC-PointNet and AC-DGCNN (see Sec. 3.2 and App. E.2). Following Esteves et al. [2018], Deng et al. [2021], we perform on-the-fly rotation augmentation during training, where the dataset size remains unchanged, and for the test set, each test example is arbitrarily rotated. We compare with PointNet, DGCNN, equivariant networks VN-PointNet and VN-DGCNN from Deng et al. [2021], canonicalization methods CN-PointNet and CN-DGCNN from Kaba et al. [2023], and traditional augmentation baselines where the training set is statically expanded with pre-generated rotated samples, denoted PointNet-Aug and DGCNN-Aug. For further details, see App. F. Tab. 3 shows the classification performance. We see our method outperforms both equivariant models and canonicalization baselines, as it allows the model to learn both local geometric features and optimal reference alignments for each class.
5 Conclusions
We introduce adaptive canonicalization based on prior maximization, a general framework for equivariant machine learning in which the standard form depends on both the input and the network. We prove that our method is continuous, symmetry preserving, and has universal approximation properties. We demonstrate the applicability of our theory in two settings: resolving eigenbasis ambiguities in spectral graph neural networks, and handling rotational symmetries in point clouds.
Limitations and Future Work. Our framework is naturally suited to classification tasks, and at the current scope of the paper, we did not address regression tasks. We will extend the adaptive canonicalization to regression in future work. Another limitation of our approach is that prior maximization requires solving $D$ optimizations at runtime for the $D$ classes. In future work, we will reduce this to a single optimization to improve efficiency.
Acknowledgments
This research was supported by a grant from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel, and the United States National Science Foundation (NSF), (NSF-BSF, grant No. 2024660), and by the Israel Science Foundation (ISF grant No. 1937/23).
Appendix
This supplemental material is structured as follows:
  • Appendix A provides an extended review of related work in equivariant machine learning, including further discussions on canonicalization methods, frame-averaging, and equivariant architectures.
  • Appendix B presents additional theoretical analysis of universal approximation theorems. We discuss the universal approximation of Euclidean functions as well as multi-sets functions.
  • Appendix C includes additional examples of continuous prior maximization. Here, we present the continuous prior maximization for unbounded point clouds under rotations, for continuous to discrete images with rotations and other transformations, as well as for discrete to discrete images with rotations and other image transformations.
  • Appendix D reviews the functional calculus and demonstrates how to plug self-adjoint operators inside functions, specifically in the context of the spectral filters. In addition, we also include a discussion on the automorphism invariance of the spectral filter.
  • Appendix E describes the construction of adaptive canonicalization applied to spectral graph neural networks and point cloud networks, as discussed in Sec. 3. We introduce how to construct the anisotropic nonlinear spectral filters and how the adaptive canonicalization method is applied to PointNet and DGCNN.
  • Appendix F outlines the experimental setup and details for the experiments in Sec. 4. We report the used datasets, including the toy problem, TUDataset, and OGB benchmarks. We include the dataset split, competing methods, and hyperparameters.
  • Appendix G presents additional results of our method, including trade-offs between sampling- and optimization-based adaptive canonicalization, the node-level anisotropic spectral filters, and the generalization of point cloud under out-of-distribution rotations.
Appendix A Related Work
We provide an extended discussion of related work for equivariant machine learning.
A.1 Canonicalization Methods: Fixed and Learned Approaches
Canonicalization [Babai and Luks, 1983, Palais and Terng, 1987, Ma et al., 2023, 2024] is a classical strategy for handling symmetry in data, especially for tasks where invariance or equivariance to group actions is desirable [Gerken et al., 2023]. It preprocesses each input by mapping it to a standard form prior to downstream learning and inference, so that all symmetry-equivalent inputs are treated identically by the subsequent model. There are two common ways for canonicalization: fixed or learned approaches.
  • Fixed Canonicalization. Fixed canonicalization uses deterministic and often analytic procedures to assign a unique representative to each symmetry orbit. For example, principal component analysis (PCA) [Jolliffe and Cadima, 2016] alignment canonically orients an object (e.g., a point cloud or molecule) by rotating it so its principal components align with the coordinate axes [Kazhdan et al., 2003]. Procrustes analysis [Gower, 1975] canonically orients sets of points by finding the optimal rotation, translation, and scale that minimizes squared point-to-point distances to a reference. For sets and graphs, canonicalization can be achieved by reordering nodes, atoms, or features so that isomorphic inputs share a single labeling. In spectral methods and spectral graph neural networks where eigenvectors are fundamental [Kipf, 2016, Defferrard et al., 2016a, Von Luxburg, 2007, Belkin and Niyogi, 2003, Dwivedi et al., 2023, Maskey et al., 2022], canonicalization of spectral decomposition [Lim et al., 2022, 2023, Ma et al., 2023, 2024] addresses eigenbasis ambiguity [Chung, 1997, Spielman, 2012] by processing each eigenspace independently and selecting representative eigenvectors or directions by applying orthogonal or axis-based projections, typically as a graph preprocessing step. An alternative approach is eigenbasis canonicalization via the input signal, where the signal itself is used to define a canonical spectral representation, making the spectral transformation independent of the arbitrary choice of eigenvectors [Lin et al., 2024a, Geisler et al., 2024].
  • Learned Canonicalization. Learned canonicalization [Zhang et al., 2018, Kaba et al., 2023, Luo et al., 2022] seeks to overcome the rigidity and inflexibility of fixed rules with a trainable mapping that selects a representative for each symmetry orbit. The canonicalizer is parameterized (typically as a neural network and trained to produce canonical forms. For example, Kaba et al. [2023] developed a neural network that learns the canonicalization transformation, which enables plug-and-play equivariance, e.g., orthogonalizing learned features via the Gram-Schmidt process [Trefethen and Bau, 2022]. Their results show that the learned canonicalizers outperform fixed canonicalizers.
However, Dym et al. [2024] pointed out that regardless of whether the canonicalizations are learned or not, a continuous canonicalization does not exist for many common groups (e.g., Sn, SO(d), O(d) on point clouds $n\geq d$). Therefore, while learned canonicalization improves empirical performance, it remains generally discontinuous and can induce instability, hinder generalization, limit model reliability on boundary cases, or out-of-distribution data. In contrast, our adaptive canonicalization framework learns the optimal transformation for each input by maximizing the predictive confidence of the network, resulting in a continuous and symmetry preserving mapping.
Symmetry Adaptation for Pretrained Models via Canonicalization.
The recent work [Mondal et al., 2023] addresses symmetry adaptation in large-scale networks by learned canonicalization before input is processed by a pretrained network. Therefore, it can convert pretrained deep networks to handle symmetry, bypassing the need to redesign or retrain the network. However, the canonicalization mapping itself is not continuous, and misalignments between canonical forms and the pretrained model’s priors can still limit effectiveness. In addition, it is important to note that its canonical form is determined by the canonicalizer’s internal objective and (potentially) supplemented with dataset priors. In contrast, our adaptive canonicalization based on prior maximization adapts the canonicalization to the downstream predictor by explicitly selecting, for each input, the transformation that maximizes the predictive confidence of the prediction model for inference. Another recent extension of canonization work [Singhal et al., 2025] explores a set of transformations at test time and uses the scoring functions of large pre-trained foundation models like CLIP [Radford et al., 2021] or SAM [Kirillov et al., 2023] to select the most “canonical” representation upon which downstream inference is performed. Note that the work in Singhal et al. [2025] does not involve any network retraining, uses the foundation models as is, and performs canonicalization entirely at inference by optimizing over transformations. While it achieves strong empirical performance, its canonicalization mapping is not continuous, and therefore, small input changes could cause abrupt switches in the selected canonical view.
A.2 Frame Averaging
Frame averaging [Puny et al., 2021] achieves equivariance to group symmetries by averaging a network’s output over a set of group transformations (known as a “frame”). It is built on the classical group averaging operator, which guarantees symmetries by summing a function over all group elements. Frame averaging has two main advantages: 1) it allows adaptation of standard non-equivalent network architectures to handle symmetry, similar to canonicalization methods, and 2) it avoids computational intractability of full group averaging, especially for large or continuous groups. Recent work [Lin et al., 2024b] proposes minimal frame averaging that attains strong symmetry coverage with small frames. Domain-specific frame averaging methods [Duval et al., 2023b, Atzmon et al., 2022] show that it can be deployed in material modeling and geometric shape analysis. However, it requires a careful selection of a suitable frame. In addition, frame averaging uses a fixed set of transformations independent of the input or task, potentially leading to sub-optimal or less discriminative feature representations. In contrast, our adaptive canonicalization learns the optimal transformation for each input in a data- and network-dependent way, yielding symmetry preserving continuous functions that can improve representation quality and empirical task performance.
A.3 Equivariant Architectures
Equivariant architectures are a class of models explicitly designed to respect symmetry groups acting on the data. Formally, a network is equivariant to a group of transformations if, when the input is transformed by a symmetry group action, the output transforms via the same group action. That is, for a group $G$ and a function $f$, equivariance guarantees $f(g\cdot x)=g\cdot f(x)$ for all $g\in G$ and input $x$. It has been developed across images [Cohen and Welling, 2016a, b, Kondor and Trivedi, 2018, Worrall et al., 2017, Weiler et al., 2018b], graphs [Bronstein et al., 2017, Zaheer et al., 2017, Gilmer et al., 2017, Maron et al., 2019, Kofinas et al., 2024, Thiede et al., 2020, Vignac et al., 2020, Keriven and Peyré, 2019], molecules [Thomas et al., 2018, Brandstetter et al., 2021, Anderson et al., 2019, Fuchs et al., 2020, Satorras et al., 2021, Duval et al., 2023a, Schütt et al., 2017, Liao and Smidt, 2023, Hordan et al., 2025, Du et al., 2022, Passaro and Zitnick, 2023], and manifolds [Masci et al., 2015, Monti et al., 2017, Cohen et al., 2019, Weiler et al., 2021]. Notably, equivariant networks often demand group-specific designs that rely on group theory, representation theory, and tensor algebra. This can reduce flexibility and raise compute and memory costs [Pertigkiozoglou et al., 2024, Wang et al., 2023, Liao and Smidt, 2023]. Making nonlinearities, pooling, and attention strictly equivariant further constrains layer choices and can increase parameter count and runtime. Moreover, imposing symmetry throughout the stack may limit the expressivity when data only approximately respect the assumed symmetry or contain symmetry-breaking noise [Wang et al., 2022, Lawrence et al., 2025]. In contrast, our approach handles symmetry by learning an input- and task-dependent canonical form through prior maximization and applying a standard backbone. By construction, this mapping is continuous and symmetry preserving. Our adaptive canonicalization removes heavy group-specific layers, reduces per-layer equivariant cost, and keeps ordinary nonlinearities and pooling.
Appendix B Universal Approximation Theorems
B.1 Universal Approximation of Euclidean Fucntions
Here, we cite a classical result stating that multilayer perceptrons (MLPs) are universal approximators of functions over compact sets in Euclidean spaces.
Theorem 14 (Universal Approximation Theorem [Leshno et al., 1993]).
Let $\sigma:{\mathbb{R}}\to{\mathbb{R}}$ be a continuous, non-polynomial function. Then, for every $M,L\in{\mathbb{N}}$, compact ${\mathcal{K}}\subseteq{\mathbb{R}}^{M}$, continuous function $f:{\mathcal{K}}\to{\mathbb{R}}^{L}$ , and $\varepsilon>0$, there exist $D\in{\mathbb{N}},{\bm{W}}_{1}\in{\mathbb{R}}^{D\times M},{\bm{b}}_{1}\in{\mathbb{R}}^{D}$ and ${\bm{W}}_{2}\in{\mathbb{R}}^{L\times D}$ s.t.
$\sup_{{\bm{x}}\in\mathcal{K}}\;\left\lvert f({\bm{x}})-{\bm{W}}_{2}\sigma\left({\bm{W}}_{1}{\bm{x}}+{\bm{b}}_{1}\right)\right\rvert\leq\varepsilon.$ (4)
B.2 Universal Approximation of Multi-Sets Functions
Multisets are sets where repetitions of elements are allowed. Formally, a multiset of $N$ elements in ${\mathbb{R}}^{J}$ can be defined as a set of pairs $(x,i)$ where $x$ denotes a point and $i$ the number of times the point $x$ appears in the multiset.
Standard universal approximation analysis of multi-set functions goes along the following lines. First, we represent multisets as 2D arrays ${\mathbf{X}}\in{\mathbb{R}}^{N\times J}$, where each row ${\mathbf{X}}_{n,:}\in{\mathbb{R}}^{J}$ represents one point in the multiset. Note that the same multi-set can be represented by many arrays. In fact, two arrays $\mathbf{X}$} and $\mathbf{X}$$^{\prime}$ represent the same multi-set if and only if one is a permutation of the other.
To formulate this property, let $\mathcal{S}_{N}$ be the symmetric group of $N$ elements, i.e., the group of permutations of $N$ elements. Given $s\in\mathcal{S}$ and ${\mathbf{X}}\in{\mathbb{R}}^{N\times J}$, let $\rho(s){\mathbf{X}}$ denote the permutation of $\mathbf{X}$} via $s$. By convention, permutations change the order of the $N$ rows of $\mathbf{X}$}, and keeps each row intact. Now, $\mathbf{X}$} and $\mathbf{X}$$^{\prime}$ represent the same multi-set if and only if there is a permutation $s\in\mathcal{S}_{N}$ such that ${\mathbf{X}}=\rho(s){\mathbf{X}}^{\prime}$.
Now, for a function $y:{\mathbb{R}}^{N\times J}\rightarrow{\mathbb{R}}^{D}$ to represent a multi-set function, it should be invariant to permutations, i.e., for every $s\in\mathcal{S}_{N}$ we have $y(\rho(s){\mathbf{X}})=y({\mathbf{X}})$. Hence, standard UATs of multi-set functions are formulated based on the following notion of universal approximation.
Definition 15.
Let $\mathcal{K}\subset{\mathbb{R}}^{N\times J}$ be an invariant compact domain, i.e., $\rho(s){\mathbf{X}}\in\mathcal{K}$ for every ${\mathbf{X}}\in\mathcal{K}$ and $s\in\mathcal{S}_{N}$. A set of invariant functions $\mathcal{N}(\mathcal{K},{\mathbb{R}}^{D})\subset C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ is called an invariant universal approximator of $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ equivariant functions if for every invariant function $y\in C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ and $\epsilon>0$ where is $\theta\in\mathcal{N}(\mathcal{K},{\mathbb{R}}^{D})$ such that for every ${\mathbf{X}}\in\mathcal{K}$
$\left\lvert\theta({\mathbf{X}})-y({\mathbf{X}})\right\rvert<\epsilon.$
For example, DeeptSets are universal approximators of invariant $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ functions [Wagstaff et al., 2022].
In this section, we describe an alternative, but equivalent, approach to model multisets of size $N$ and their universal approximation theorems using the notion of quotient. The motivation is that our main UAT, Theorem 13, is based on the standard symmetryless notion of universal approximation, Definition 3. While it is possible to develop our adaptive canonicalization theory for functions that preserve symmetries, and obtain an analogous theorem to Theorem 13 based on invariant universal approximators, there is no need for such complications. Instead, we can use the standard definition of universal approximation (Definition 3), and directly encode the symmetries in the domain using quotient spaces, as we develop next.
Quotient Spaces.
Let $\mathcal{X}$ be a topological space, and $x\sim y$ an equivalence relation between pairs of points. The equivalence class $[x]$ of $x\in\mathcal{X}$ is defined to be the set
$x]:=\{y\in X\ x\sim y\}.$
Definition 16 (Quotient topology).
Let $\mathcal{X}$ be a topological space, and $\sim$ an equivalence relation on $\mathcal{X}$. The quotient set is defined to be
$\mathcal{X}/\sim:=\{[x]\ |\ x\in\mathcal{X}\}.$
The quotient set is endowed with the quotient topology. The quotient topology is the finest (largest) topology making the mapping $\nu:x\mapsto[x]$ continuous. In other words, the open sets $B\subset(\mathcal{X}/\sim)$ are those sets such that $\cup_{[x]\in B}[x]$ is open in $\mathcal{X}$.
The mapping $\nu:\mathcal{X}\rightarrow(\mathcal{X}/\sim)$, defined by $\nu(x)=[x]$, is called the canonical projection.
Multi-Sets as Equivalence Classes and UATs.
Define the equivalence relation: ${\mathbf{X}}\sim{\mathbf{Y}}$ if there exists $s\in\mathcal{S}_{N}$ such that ${\mathbf{X}}=\rho(s){\mathbf{Y}}$. Now, a multi-set of $N$ elements can be defined as ${\mathbb{R}}^{N\times J}/\sim$. As opposed to the definition of multisets as sets of pairs $(x,i)$, the quotient definition automatically gives a topology to the sets of multisets, namely, the quotient topology. In fact, it can be shown that the quotient topology is induced by the following metric.
Definition 17 (Multi-Set Metric).
Given ${\mathbf{X}}],[{\mathbf{Y}}]\in(\mathbb{R}^{N\times J}/\sim)$, their distance is defined to be
$d([{\mathbf{X}}],[{\mathbf{Y}}]):=\min_{{\mathbf{X}}^{\prime}\in[{\mathbf{X}}],{\mathbf{Y}}^{\prime}\in[{\mathbf{Y}}]}\|{\mathbf{X}}^{\prime}-{\mathbf{Y}}^{\prime}\|.$
It is easy to see that
$d([{\mathbf{X}}],[{\mathbf{Y}}]):=\min_{s\in\mathcal{S}_{N}}\|{\mathbf{X}}-\rho(s){\mathbf{Y}}\|.$
One can show that the above distance is indeed a metric, and that the topology induced by this metric is exactly the quotient topology, i.e., $d$ metrizes $({\mathbb{R}}^{N\times J}/\sim)$.
Theorem 18.
The metric $d([{\mathbf{X}}],[{\mathbf{Y}}])$ metrizes the quotient topology ${\mathbb{R}}^{N\times J}/\sim$.
A set $\mathcal{K}\subset{\mathbb{R}}^{N\times J}$ is called invariant if for every ${\mathbf{X}}\in\mathcal{K}$ and $s\in\mathcal{S_{N}}$ we have $\rho(s){\mathbf{X}}\in\mathcal{K}$. Consider the quotient space
$\mathcal{K}/\sim=\{[{\mathbf{X}}]\ |\ {\mathbf{X}}\in\mathcal{K}\}\subset{\mathbb{R}}^{N\times J}/\sim.$
We now have the following proposition about continuity of symmetric functions.
Proposition 19.
Let $\mathcal{K}\subset{\mathbb{R}}^{N\times J}$ be an invariant compact domain. For every continuous invariant mapping $y:\mathcal{K}\rightarrow{\mathbb{R}}^{D}$ there exists a unique continuous mapping $\overline{y}:(\mathcal{K}/\sim)\rightarrow{\mathbb{R}}^{D}$ such that
$y=\overline{y}\circ\nu,$
where $\nu$ is the canonical projection. On the other hand, for every function $z\in C_{0}(\mathcal{K}/\sim,{\mathbb{R}}^{D})$, we have that $z\circ\nu$ is a continuous invariant function in $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$.
Let $\mathcal{K}\subset{\mathbb{R}}^{N\times j}$ be an invariant domain. For a set of continuous invariant functions $\mathcal{N}\subset C_{0}(\mathcal{K},{\mathbb{R}}^{D})$, we denote
$(\mathcal{N}/\sim):=\{\overline{y}\ |\ y\in\mathcal{N}\}\subset C_{0}(\mathcal{K}/\sim,{\mathbb{R}}^{D}).$
Note that by Proposition 19
$\big(C_{0}(\mathcal{K},{\mathbb{R}}^{D})/\sim\big)=C_{0}\big(\mathcal{K}/\sim,{\mathbb{R}}^{D}\big).$
This immediately leads to a UAT theorem for multi-set continuous functions in which every continuous multi-set function can be approximated by a neural network.
Theorem 20.
Let $\mathcal{K}\subset{\mathbb{R}}^{N}$ be an invariant compact domain, and let $\mathcal{N}(\mathcal{K},{\mathbb{R}}^{D})\subset C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ be an invariant universal approximator of $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ equivariant functions. Then $\mathcal{N}(\mathcal{K},{\mathbb{R}}^{D})/\sim$ is a universal approximator $C_{0}(\mathcal{K}/\sim,{\mathbb{R}}^{D})$.
Note that in the above UAT the symmetries are directly encoded in the quotient spaces, and, hence, there is no need to encode any symmetry in the spaces of functions. Hence, Theorem 20 is based on the standard symmetryless definition of universal approximation – Definition 3 – rather than the symmetry driven construction of Definition 15. As a result, we can directly use our theory of adaptive canonicalization on multi-set functions. Specifically, we can use Theorem 13, where the space $\mathcal{K}$ in the theorem is taken as $\mathcal{K}/\sim$ in our above analysis.
Now, we immediately obtain that the set of neural networks $\overline{\theta}$ where $\theta$ is a DeepSet is a universal approximator of the space of continuous multi-set functions.
Appendix C Additional Examples of Continuous Prior Maximization
We first note that when $\mathcal{K}$ is a locally compact metric space, functions in $C_{0}(\mathcal{K},{\mathbb{R}}^{D})$ must be uniformly continuous.
Unbounded Point Clouds and Rotations.
Let $\mathcal{U}=\mathcal{SO}(3)$ be the space of 3D rotations, and $\mathcal{G}=\mathcal{K}={\mathbb{R}}^{N\times 3}$ the set of sequences of $N$ points in ${\mathbb{R}}^{3}$, i.e. the space of point clouds. Consider the $\mathcal{L}_{2}$ metric in $\mathcal{G}$. Consider the rotation $g\mapsto\kappa_{u}(g)$ of the point cloud $g$ by $u\in\mathcal{U}$.
Let $f\in C_{0}(\mathcal{G},{\mathbb{R}}^{D})$. Next we show that $f$ must be uniformly continuous. Let $\epsilon>0$. By the fact that $f$ vanishes at infinity, there exists a compact domain $\mathcal{K}\subset{\mathbb{R}}^{N\times 3}$ such that for every $x\notin\mathcal{K}$ we have $\left\lvert f(x)\right\rvert<\epsilon/2$. By the fact that $\mathcal{U}\times\mathcal{K}$ is compact and $\kappa$ continuous, $\kappa$ is uniformly continuous on $\mathcal{U}\times\mathcal{K}$. Hence, there exists $\delta_{\epsilon}>0$ such that every $g,g^{\prime}\in\mathcal{K}$ with $d(g,g^{\prime})<\delta_{\epsilon}$ satisfy $d(f(g),f(g^{\prime}))<\epsilon$. Let $\kappa^{\prime}$ be the compact space consisting of all point of distance less or equal to $\delta_{\epsilon}$ from $\mathcal{K}$. There exists $0<\delta^{\prime}_{\epsilon}<\delta_{\epsilon}$ such that every $g,g^{\prime}\in\mathcal{K}^{\prime}$ with $d(g,g^{\prime})<\delta^{\prime}_{\epsilon}$ satisfy $d(f(g),f(g^{\prime}))<\epsilon$.
Now, let $g,g^{\prime}\in\mathcal{G}$ satisfy $d(g,g^{\prime})<\delta^{\prime}_{\epsilon}$. If one of the point $g$ or $g^{\prime}$ lies outside $\mathcal{K}^{\prime}$, then both of them lie outside $\mathcal{K}$, so
$d(f(g),f(g^{\prime}))\leq\|f(g)\|+\|f(g^{\prime})\|<\epsilon.$
Otherwise, both lie in $\mathcal{K}^{\prime}$, so $d(f(g),f(g^{\prime}))<\epsilon$. Both cases together mean that $f$ is uniformly continuous.
As a result of uniform continuity, $\{g\mapsto f(\kappa_{u}(g))\}){u\in\mathcal{U}}$ is equicontinuous, and this is a setting of continuous prior maximization.
In fact, this analysis shows that whenever $\mathcal{G}=\mathcal{K}$ and $\kappa$ is continuous in $(u,g)$, then it’s corresponding $\rho$ is continuous prior maximization.
Continuous to Discrete Images with Rotations and Other Image Transformations.
Consider the “continuous” space of images $\mathcal{G}=\mathcal{L}_{2}({\mathbb{R}}^{2})$ and the discrete space $\mathcal{K}={\mathbb{R}}^{N\times N}$ of images of $N\times N$ pixels.
Let $\mathcal{U}$ be the unit circle. For $g\in\mathcal{G}$ and $u\in\mathcal{U}$ let $\pi(u)g$ be the rotation of the image $g$ by angle $u$. To define the discretizing mapping $P:\mathcal{L}_{2}({\mathbb{R}}^{2})\rightarrow{\mathbb{R}}^{N\times N}$, consider the partition of [1,1][-1,1] into the $N$ intervals
$I_{n}=[-1+2n/N,-1+2(n+1)/N),\quad n=0,\ldots N-1.$
Consider the closed linear subspace $\mathcal{D}({\mathbb{R}}^{2})\subset\mathcal{L}_{2}({\mathbb{R}}^{2})$ consisting of images that are zero outside $1,1]^{2}$ and piecewise constant on the squares $\{I_{n}\times I_{m}\}_{n,m=0}^{N-1}$. Now, $P$ is the operator that takes $g\in\mathcal{L}_{2}({\mathbb{R}}^{2})$ first orthogonally projects it upon $\mathcal{D}({\mathbb{R}}^{2})$ to get $g^{\prime}$, and then returns
$P(g)=\{g^{\prime}(-1+2n/N,-1+2m/N)\}_{n,m=0}^{N-1}\in{\mathbb{R}}^{N\times N}.$
Define the mapping $\kappa_{u}$ as follows: $\kappa_{u}(g)=P(\pi(u)g)$. By the fact that $\pi_{u}$ is an isometry for every $u\in\mathcal{U}$ and $P$ is non-expansive (as an orthogonal projection), $\kappa_{u}:\mathcal{G}\rightarrow{\mathbb{R}}^{N\times N}$ is Lipschitz 1 for every $u$. Hence, $\{g\mapsto f(\kappa_{u}(g))\}_{u\in\mathcal{U}}$ is equicontinuous.
As a result, the corresponding prior minimization is a continuous prior minimization.
This setting can be extended to other image deformation based on diffeomorphisms of the domain ${\mathbb{R}}^{2}$, parameterized by compact spaces $\mathcal{U}$, For example, one can take dilations up to some uniformly bounded scale. More generally, One can consider a compact set $\mathcal{U}$ of matrices in ${\mathbb{R}}^{2\times 2}$, and define $\pi(u)g(\mathbf{x})=g(\mathcal{x}\mathbf{U})$ for $\mathbf{U}\in\mathcal{U}$, $g\in L_{2}({\mathbb{R}}^{2})$ and $\mathbf{x}\in{\mathbb{R}}^{2}$.
Discrete to Discrete Images with Rotations and Other Image Transformations.
One can replace $\mathcal{G}$ by $\mathcal{D}({\mathbb{R}}^{2})$ in the above analysis. Since now $\mathcal{D}({\mathbb{R}}^{2})$ is compact, $\rho$ must be a continuous prior maximization.
Appendix D Functional Calculus and Spectral Filters
In this section, we recall the theory of plugging self-adjoint operators inside functions.
Spectral Theorem.
Let $\mathcal{L}$ be a self-adjoint operator on a finite-dimensional Hilbert space (e.g., $\mathcal{L}\in\mathbb{C}^{N\times N}$ with $\mathcal{L}=\mathcal{L}^{*}$). There exists a unitary matrix $V$ and a real diagonal matrix $\Lambda=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{N})$ such that $\mathcal{L}=V\Lambda V^{*}$. The columns $v_{i}$ of $V$ form an orthonormal eigenbasis with $\mathcal{L}v_{i}=\lambda_{i}v_{i}$.
Functional Calculus.
For any function $f:\mathbb{R}\to\mathbb{C}$ defined on the spectrum $\sigma(\mathcal{L})=\{\lambda_{1},\dots,\lambda_{N}\}$:
$\quad f(\mathcal{L})\;:=\;Vf(\Lambda)V^{*},\qquad f(\Lambda)=\mathrm{diag}\big(f(\lambda_{1}),\dots,f(\lambda_{N})\big).\quad$
Equivalently, we can write the spectral projections $P_{i}:=v_{i}v_{i}^{*}$,
$f(\mathcal{L})\;=\;\sum_{i=1}^{N}f(\lambda_{i})\,P_{i}.$
We can plug a self-adjoint matrix into a function by: 1) diagonalize $\mathcal{L}$, 2) apply $f$ to the eigenvalues, and 3) conjugate back.
Take $f=\mathbbm{1}_{I}$ for a Borel set $I\subset{\mathbb{R}}$. The indicator function $\mathbbm{1}_{I}(\mathcal{L})$ is an orthogonal projection, since $\mathbbm{1}_{I}(\mathcal{L})^{2}=\mathbbm{1}_{I}(\mathcal{L})$ and $\mathbbm{1}_{I}(\mathcal{L})^{*}=\mathbbm{1}_{I}(\mathcal{L})$.
Spectral Graph Filters.
A graph shift operator (GSO) is a self-adjoint matrix that reflects the graph’s connectivity, such as a (normalized) graph Laplacian or a symmetrized adjacency. Let $\mathcal{L}$ be such a GSO with eigenpairs $\{(\lambda_{i},v_{i})\}_{i=1}^{N}$ and $V=[v_{1},\ldots,v_{N}]$. For a $T$-channel node signal $\mathbf{X}\in\mathbb{R}^{N\times T}$ and a matrix-valued frequency response $g:\mathbb{R}\to\mathbb{R}^{d^{\prime}\times T}$, the spectral filter
$g(\mathcal{L})\mathbf{X}:=\sum_{i=1}^{N}v_{i}v_{i}^{\top}\,\mathbf{X}\,g(\lambda_{i})^{\top}$
applies the graph convolution theorem [Bracewell and Kahn, 1966] with $d^{\prime}$ output channel: each Fourier mode $v_{i}$ is preserved in space, while channels are mixed by $g(\lambda_{i})$ in the spectral domain. In the scalar case ($T=d^{\prime}=1$) with $f:\mathbb{R}\to\mathbb{R}$, the spectral filter simply reduces to the functional-calculus operator acting on $\mathbf{X}$:
$f(\mathcal{L})\mathbf{X}=\sum_{i=1}^{N}f(\lambda_{i})\,v_{i}v_{i}^{\top}\mathbf{X}\;=\;Vf(\Lambda)V^{\top}\mathbf{X}.$
Spectral graph neural networks [Defferrard et al., 2016b, Kipf, 2016, Levie et al., 2018] compose such filters with pointwise nonlinearities, using trainable $g$ at each layer.
Automorphism Invariance of Spectral Filters.
By construction, a spectral filter depends only on the eigenvalues and eigenvectors of the GSO $\mathcal{L}$. Let $\pi$ be a graph automorphism (i.e., a permutation of vertices that preserves adjacency and hence commutes with $\mathcal{L}$), the operator satisfies
$P_{\pi}\mathcal{L}P_{\pi}^{\top}=\mathcal{L},$
where $P_{\pi}$ is the permutation matrix associated with $\pi$. By functional calculus, $f(\mathcal{L})=P_{\pi}f(\mathcal{L})P_{\pi}^{\top}$ holds for any spectral filter $g$. This means that the action of a spectral filter does not change under any automorphism of the graph, i.e., its output is invariant to node re-labeling.
When we apply a spectral filter on a square grid graph, the response along the $x$-axis and along the $y$-axis is the same. This implies that the spectral filters are isotropic in the spatial domain, and therefore, the spectral filters can not distinguish between directions, such as horizontal vs. vertical, and impose directional limitations.
Appendix E Application of Adaptive Canonicalization to Anisotropic Geometric Networks: Construction Details
In this section, we present the construction details of the application of adaptive canonicalization to anisotropic geometric networks
E.1 Construction Details for Anisotropic Nonlinear Spectral Filters
We introduce anisotropic nonlinear spectral filters as a mechanism to perform adaptive canonicalization for eigenbasis selection within the framework of spectral filtering. Unlike standard approaches, which apply uniform (isotropic) filtering across the spectrum, our formulation partitions the spectrum into disjoint frequency bands and treats each independently. Within every band, we introduce a unitary transformation that reorients the eigenspace based on prior maximization, thereby resolving orientation ambiguities inherent in spectral methods.
In the spectral filter, we consider the GSO as the normalized graph Laplacian $\mathcal{L}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, where $\mathbf{A}$ is the adjacency matrix and $\mathbf{D}$ is a diagonal degree matrix with entries ${\mathbf{d}}=(d_{i})_{i}\in\mathbb{R}^{N}$. The eigendecomposition of $\mathcal{L}$ is given by $\mathcal{L}={\mathbf{V}}\mathbf{\Lambda}{\mathbf{V}}^{\top}$, where ${\mathbf{V}}=({\mathbf{v}}_{i})_{i}\in{\mathbb{R}}^{N\times N}$ is the matrix of eigenvectors (i.e., the eigenbasis) and $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{N})$ is the diagonal matrix of eigenvalues, where $\lambda_{1}\leq\ldots\leq\lambda_{N}$. Consider predefined bands $b_{0}<b_{1}<\ldots<b_{B}$ where $b_{0},b_{B}]\subseteq[0,2]$. In particular, we adopt a dyadic partitioning scheme where the subbands are defined as $\left[\lambda_{N}r^{S-k+1},\lambda_{N}r^{S-k}\right]$ with $0<r<1$ denoting the decay rate, following the approach introduced in Lin et al. [2024b]. We note that band design can be realized through different mechanisms, such as an attention mechanism, e.g., SpecFormer [Bo et al., 2023]. We leave this exploration for future work.
The indicator functions $P_{k}:=\mathbbm{1}_{[b_{k-1},b_{k})}:{\mathbb{R}}\rightarrow{\mathbb{R}}_{+}$ are the orthogonal projections onto the eigenspace corresponding to the $k$-th band. We augment these projections with band-wise unitary actions that commute with the GSO $\mathcal{L}$. Specifically, for each spectral band $k$ we consider a unitary matrix ${\mathbf{U}}_{k}\in\mathcal{O}(M_{k})$, where $M_{k}=\mathrm{rank}(P_{k})$. The global transformation operator is thus block-diagonal in the spectral domain ${\mathbf{U}}=\sum_{k}{\mathbf{U}}_{k}P_{k}\equiv\bigoplus_{j}{\mathbf{U}}_{k}$. Each eigenspace is treated independently but with a learnable, band-specific orientation.
We consider a classification task with $D$ classes and deploy the prior maximization as follows. Specifically, we introduce $D$ class-specific priors and orientations. For each class $d\in\{1,\ldots,D\}$ and band $k$, let $\mathbf{U}^{(d)}_{k}\in\mathcal{O}(M_{k})$ and define the block operator $\mathbf{U}^{(d)}\!\!=\bigoplus_{k=1}^{B}\mathbf{U}^{(d)}_{k}$. Let ${\mathbf{X}}\in{\mathbb{R}}^{N\times T}$ be the input with $T$ channels per node and let ${\mathbf{V}}_{k}\in\mathbb{R}^{N\times M_{k}}$ be the matrix of eigenvectors spanning the $k$-th band. The class-$d$ band coefficients are
$\widetilde{\mathbf{X}}^{(d)}_{k}=\mathbf{U}^{(d)}_{k}\,\mathbf{V}_{k}^{\top}\mathbf{X}\in\mathbb{R}^{M_{k}\times T}.$
Each class has its own prior that selects an orientation that stresses features for that class. We denote $h_{d}$ the class-$d$ prior. The class-$d$ orientation is obtained by
$\mathbf{U}^{\square(d)}=\bigoplus_{k}{\mathbf{U}}_{k}^{\square(d)}=\underset{\mathbf{U}^{(d)}_{1},\ldots,\mathbf{U}^{(d)}_{B}}{\arg\max}\;h_{d}(\Psi_{d}(\phi(\widetilde{\mathbf{X}}^{(d)}_{1},\ldots,\widetilde{\mathbf{X}}^{(d)}_{B})))$,
where $\phi$ is a mapping on band-wise features, and $\Psi_{d}$ is the class-specific classifier for class $d$. The nonlinear aspect emerges from the mapping network on the spectral coefficients, which operates on the band-wise features. The optimal orientations are then applied band-wise and class-wise, and the network output is computed with respect to this canonical eigenbasis. The anisotropy stems from the independent reorientation of spectral bands, in contrast to isotropic formulations that enforce a single global alignment of the spectrum.
One vs. Rest (Independent Binary Classifiers).
In each class-specific classifier, we apply a scoring function to the prior’s output. Specifically, we consider the one vs. rest setup [Rifkin and Klautau, 2004, Galar et al., 2011, Allwein et al., 2000]. We obtain the class logit
$s_{d}=\Psi_{d}(\phi(\widetilde{\mathbf{X}}^{\square(d)}_{1},\ldots,\widetilde{\mathbf{X}}^{\square(d)}_{B})),\quad\widetilde{\mathbf{X}}^{\square(d)}_{k}:=\mathbf{U}^{\square(d)}_{k}\,\mathbf{V}_{k}^{\top}\mathbf{X}.$
We then define the per-class score by a sigmoid $\hat{y}_{d}=\sigma(s_{d})$. Note that the vector $(\hat{y}_{1},\ldots,\hat{y}_{D})$ is not a probability measure, since each entry represents the independent probability of class $c$ being present. We use binary cross-entropy per class and sum over classes:
$\sum_{d=1}^{D}(-\,y_{d}\log\sigma(s_{d})-(1-y_{d})\log(1-\sigma(s_{d}))).$
Note that we can include class weights $w_{d}\geq 0$ to handle imbalance between classes. Then, at test time we predict the class with the largest logit $\arg\max_{d}s_{d}$.
We develop the graph-level representation of adaptive canonicalization applied to spectral methods, which is fully spectral. Details on the corresponding node-level representation are provided in App. G.2.
Optimizing Unitary Orientations for Band-Specific Representation.
For multiple graphs, $M_{k}$ may vary; to standardize dimensions, each MLP $\phi$ uses the largest band size and smaller bands are zero-padded. From an implementation standpoint, we can search over unitary orientations $\{{\mathbf{U}}_{k}^{(\cdot)}\}_{k=1}^{B}$, one per band, and select the pooled set of orientations that maximizes the prior coupled with the classification task. One option is to generate a set of Haar-distributed random unitary matrices for each band [Mezzadri, 2006]. Based on the chosen prior objective, the network then selects the orientation to retain. This procedure, sampling random unitaries and picking the one that maximizes the prior, should be viewed as one possible implementation rather than a requirement. Alternatively, the prior can be maximized more effectively by refining the sampled matrices. For example, we can initialize with random unitary matrices, apply the prior maximization, and then perform a few additional gradient descent (GD) steps to reduce the objective locally. In this case, the optimization is treated as an inner loop: the refined orthogonal matrix is plugged into the outer computation, without differentiating through the iterative gradient descent itself. Thus, gradient descent is only used as a local refinement step, not as a meta-learning procedure. This two‐stage procedure combines cheap random exploration with targeted local search to maximize the prior effectively without incurring the high cost.
Remarks. (i) The specific prior maximization run in parallel; complexity scales linearly with $D$. (ii) Sampling-and-refine (Haar initialization with a few local GD steps) applies per class and per band. (iii) A light regularizer can discourage degenerate reuse of the same orientation across classes, e.g., $\sum_{k}\sum_{d\neq d^{\prime}}\|\mathbf{U}^{\square(d)}_{k}-\mathbf{U}^{\square(d^{\prime})}_{k}\|_{F}^{2}$ when desired.
E.2 Construction Details for Anisotropic Point Cloud Networks
We next consider adaptive canonicalization in the spatial domain for points embedded in ${\mathbb{R}}^{3}$ with xyzxyz coordinates, such as molecular graphs [Duval et al., 2023a] or point clouds [Luo et al., 2022], where each node is associated with a 3D position vector. In this paper, we specifically focus on point cloud networks applied to classification tasks.
Point cloud networks operate directly on xyzxyz coordinates, creating an inherent anisotropy problem [Wiersma et al., 2022, Qian et al., 2021]. The models treat axis-labeled xyzxyz coordinates and their offsets as separate components, with linear layers processing each axis independently. However, under 3D rotation, these coordinate components mix, and since generic weight matrices do not commute with rotation, the model’s response becomes dependent on orientation [Shen et al., 2021].
To address this rotational ambiguity in point cloud classification tasks, we introduce an adaptive canonicalization module that searches over group actions in $\mathcal{SO}(3)$, the 3D rotation group. We now present the application of adaptive canonicalization (AC) to three widely used point cloud learning architectures: DeepSet [Zaheer et al., 2017], PointNet [Qi et al., 2017], and DGCNN [Wang et al., 2019], and denote them respectively AC-DeepSet, AC-PointNet, and AC-DGCNN.
Adaptive Canonicalization Applied to DeepSet (AC-DeepSet).
We first describe adaptive canonicalization for the DeepSet architecture [Zaheer et al., 2017], which is a general framework for learning function on sets and has UAT [Wagstaff et al., 2022]. Given a point cloud represented as $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\subset{\mathbb{R}}^{3}$, DeepSet computes a permutation-invariant representation by encoding each point with an MLP $\phi$, aggregating features with summation, and applying a global function $\xi$:
$\mathbf{x}^{\prime}=\xi\left(\sum_{i=1}^{N}\phi(\mathbf{x}_{i})\right)$
For classification into $D$ classes, we assign each class $d\in\{1,\ldots,D\}$ a learned canonical form ${\mathbf{R}}_{d}^{\square}\in\mathcal{SO}(3)$ for a prior $h_{d}$ chosen to maximize class separability in one vs. rest manner:
${\mathbf{R}}_{d}^{\square}=\underset{{\mathbf{R}}\in\mathcal{SO}(3)}{\arg\max}\;h_{d}\left(\Psi_{d}\left(\xi\left(\sum_{i=1}^{N}\phi(\mathbf{x}_{i}{\mathbf{R}}^{\top})\right)\right)\right),$
where $\Psi_{d}$ is a class-specific classifier for class $d$. Then, the class scores are computed by $s_{d}=\psi_{d}\left(\xi\left(\sum_{i=1}^{N}\phi(\mathbf{x}_{i}{\mathbf{R}}_{d}^{\square\top})\right)\right)$ for all $d\in\{1,\ldots,D\}$ , and then passed through sigmoids. The loss is trained by summing over $D$ binary cross-entropies.
Adaptive Canonicalization Applied to PointNet (AC-PointNet).
Next, we apply adaptive canonicalization in PointNet [Qi et al., 2017]. For a point cloud $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\in{\mathbb{R}}^{3}$, PointNet processes each point independently with a shared MLP $\phi$, and then aggregates points features using a symmetric function (e.g., max-pooling)
$\mathbf{x}^{\prime}=\operatorname{Pool}(\phi(\mathbf{x}_{1}),\ldots,\phi(\mathbf{x}_{N})).$
For each class $d$ with a prior $h_{d}$, we define a class-specific orientation ${\mathbf{R}}_{d}^{\square}$ obtained by maximizing discriminative power:
${\mathbf{R}}_{d}^{\square}=\underset{{\mathbf{R}}\in\mathcal{SO}(3)}{\arg\max}\;h_{d}(\Psi_{d}(\operatorname{Pool}(\phi(\mathbf{x}_{1}{\mathbf{R}}^{\top}),\ldots,\phi(\mathbf{x}_{N}{\mathbf{R}}^{\top})))),$
where $\Psi_{d}$ is the class-specific classifier conducting binary classification in the one vs. rest manner. Each ${\mathbf{R}}_{d}^{\square}$ emphasizes the characteristics most relevant to class $d$. The scoring head is denoted by $s_{d}=\Psi_{d}(\operatorname{Pool}(\phi(\mathbf{x}_{1}{\mathbf{R}}_{d}^{\square\top}),\ldots,\phi(\mathbf{x}_{N}{\mathbf{R}}_{d}^{\square\top})))$. Then we define the per-class head score by a sigmoid and use binary cross-entropy per class and sum over classes.
Adaptive Canonicalization Applied to DGCNN (AC-DGCNN).
We then consider the application of adaptive canonicalization to DGCNN [Wang et al., 2019]. Consider the input ${\mathbf{X}}\in{\mathbb{R}}^{N\times 3}$ as a point cloud with points $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}\in{\mathbb{R}}^{3}$. DGCNN constructs dynamic $k$-nearest graphs by computing $G=(V,E)$ where $E=\{(i,j):j\in\operatorname{kNN}(\mathbf{x}_{i},k)\}$. Then, the edge convolution is performed by computing the edge features and applying a max pooling:
$\mathbf{x}_{i}^{\prime}=\operatorname{Pool}_{(i,j)\in E}(\operatorname{ReLU}(\Psi(\mathbf{x}_{j}-\mathbf{x}_{i},\mathbf{x}_{i}))).$
Consider a classification task with $D$ classes. For every class $d\in\{1,\ldots,D\}$, we define a class-specific orientation ${\mathbf{R}}_{d}^{\square}$ obtained by optimizing a class-dependent prior $h_{d}$:
${\mathbf{R}}_{d}^{\square}=\underset{{\mathbf{R}}\in\mathcal{SO}(3)}{\arg\max}\;h_{d}(\Psi_{d}(\operatorname{Pool}_{(i,j)\in E}(\operatorname{ReLU}(\Psi((\mathbf{x}_{j}-\mathbf{x}_{i}){\mathbf{R}}^{\top},\mathbf{x}_{i}{\mathbf{R}}^{\top}))))),$ (5)
where $\Psi_{d}$ is the class-specific classifier. The canonical form ${\mathbf{R}}_{d}^{\square}$ emphasizes the characteristics most relevant to class $d$. Then, the classifier is applied to the output under ${\mathbf{R}}_{d}^{\square}$, i.e., $s_{d}=\Psi_{d}(\operatorname{Pool}_{(i,j)\in E}(\operatorname{ReLU}(\Psi((\mathbf{x}_{j}-\mathbf{x}_{i}){\mathbf{R}}^{\square\top},\mathbf{x}_{i}{\mathbf{R}}^{\square\top}))))$ and defines the per-class head score by a sigmoid. We then use binary cross-entropy per class and sum over classes.
Efficient Optimization with Random Sampling and Local Refinement.
A direct search over rotations can be implemented with quaternion [Shoemake, 1992] or Rodrigues’ formula [Rodrigues, 1840], but exhaustive exploration of all rotations is costly. For computation efficiency, we adopt a sample and refine scheme. Specifically, to optimize each class’s canonical rotation efficiently, we first sample a small set of random rotation matrices to serve as initial candidates. For each candidate, we compute the prior-weighted objective and then perform the refinement by taking the gradient descent steps. Note that this inner loop of gradient descent is not differentiated through: the refined rotation is treated as a fixed input to the subsequent forward pass and backpropagation. In this way, we avoid the high cost of full meta-optimization over all rotations. This keeps the expensive optimization out of backpropagation while providing a good canonical orientation per class.
Adaptive canonicalization operates like a max-pooling step: the network evaluates multiple candidate orientations and selects the one that maximizes the prior. The chosen orientation is passed forward, and during backpropagation, gradients flow only through this selected orientation.
Appendix F Experimental Details
In this section, we describe the experimental setups and implementation details used in Sec. 4.
F.1 Illustrative Toy Problems: Grid Signal Orientation Tasks
Toy Problem and Experimental Setup.
We consider square grid on the torus with a 2-channel signal. The first channel contains a sinusoidal signal aligned with the $x$-axis, given by $\sin(2\pi x/T)$. The second channel depends on the class label: in Class 0 it is aligned with the $x$-axis, while in Class 1 it encodes a sinusoidal signal along the $y$-axis, $\sin(2\pi y/T)$. Independent Gaussian noise with variance $\sigma^{2}$ is added to each channel. In addition, it introduces an additional challenge by spatially restricting the support of the channels. The grid is vertically partitioned into two disjoint halves. The first channel is supported only on the left half. The second channel is supported only on the right half. The task is to decide if the frequency at the two channels is in the same orientation. The grid size is fixed at $N=40^{2}$, the sinusoidal period is set to $T=20$, and the noise level is chosen as $\sigma=0.1$. We generate 1000 samples. Evaluation is carried out using 10-fold cross-validation.
Competing Methods.
The competing methods include: MLP, GCN, GAT, GIN, ChebNet, NLSF, S2GNN, FA+GIN, and OAP+GIN.
Hyperparameters.
We use a three-layer network with a hidden feature dimension chosen from $\{32,64,128\}$ and ReLU activation functions. The learning rate is selected from $\{10^{-3},10^{-4},10^{-5}\}$. Batch size 100. All models are implemented in PyTorch and optimized with the Adam optimizer [Kingma and Ba, 2014]. Experiments are conducted on an Nvidia DGX A100. The output of the GNN is then passed to an MLP, followed by a softmax classifier.
F.2 Graph Classification on TUDataset
Datasets and Experimental Setup.
We consider five graph classification benchmarks from TUDataset [Morris et al., 2020]: MUTAG, PTC, ENZYMES, PROTEINS, and NCI1. The dataset statistic is reported in Tab. LABEL:tab:statistics. Following the random split protocol [Ma et al., 2019, Ying et al., 2018, Zhang et al., 2019b], we partition the dataset into 80% training, 10% validation, and 10% testing. Results are averaged over 10 random splits, with mean accuracy and standard deviation reported.
Competing Baselines.
We evaluate on medium-scale graph classification benchmarks from TUDataset, using the same set of competing methods as in grid signal orientation tasks. The baselines include MLP, GCN, GAT, GIN, ChebNet, NLSF, S2GNN, FA+GIN, and OAP+GIN.
Hyperparameters.
The hidden dimension is set to be 128. The models are implemented using PyTorch, optimized with the Adam optimizer [Kingma and Ba, 2014]. An early stopping strategy is applied, where training halts if the validation loss does not improve for 100 consecutive epochs. The hyperparameters are selected through a grid search, conducted via Optuna [Akiba et al., 2019], with with the learning rate and weight decay explored in the set $\{1e^{-2},1e^{-3},1e^{-4}\}$, the pooling ratio varying within $[0.1,0.9]$ with step $0.1$, and the number of layers ranging from 2 to 9 in a step size of 1. The output representations are then passed into an MLP followed by a softmax layer, and predictions are obtained by optimizing a cross-entropy loss function. Experiments are conducted on an Nvidia DGX A100.
Table 4: Datasets statistics.
Dataset #\# Graphs #\# Classes Avg.#\# Nodes Avg.#\# Edges
MUTAG 188 2 17.93 19.79
PTC 344 2 14.29 14.69
ENZYMES 600 6 32.63 64.14
PROTEINS 1113 2 39.06 72.82
NCI1 4110 2 29.87 32.30
ogbg-molhiv 41127 2 25.5 27.5
ogbg-molpcba 437929 128 26.0 28.1
ogbg-ppa 158100 37 243.4 2266.1
F.3 Molecular Classification on OGB Datasets
Datasets and Experimental Setup.
We evaluate on larger-scale benchmarks from the Open Graph Benchmark (OGB) dataset [Hu et al., 2020] for classification tasks, including ogbg-molhiv, ogbg-molpcba, and ogbg-ppa. Dataset statistics are summarized in Tab. LABEL:tab:statistics The evaluation settings are followed by the OGB protocol [Hu et al., 2020].
Competing Baselines.
For large-scale graph classification, we include GCN, GIN, GatedGCN, PNA, GraphTrans, SAT, GPS, SAN, and the canonicalization-based variant OAP+GatedGCN as the competing methods. These approaches have previously demonstrated strong performance on OGB benchmarks, and their reported results are taken from prior work [^2] https://ogb.stanford.edu/docs/leader_graphprop/.
Hyperparameters.
The models are implemented in PyTorch and optimized with the Adam optimizer, with training capped at a maximum of 1000 epochs and controlled by an early stopping criterion. The hidden dimension is selected from the set $\{128,256,512\}$, while the number of layers varies from 2 to 10 in steps of 1. Dropout rates are explored within the range $0,0.1,\dots,0.5]$, the learning rate is tuned within the interval $[0.0001,0.001]$, and the warmup is set as 5 or 10. Additionally, the batch size is chosen from $\{32,64,128,256\}$ and the weight decay is chosen from $\{10^{-4},10^{-5},10^{-6}\}$. All hyperparameters are tuned using Optuna [Akiba et al., 2019]. The experiments are conducted on an NVIDIA A100 GPU.
F.4 ModelNet40 Point Cloud Classification
Datasets and Experimental Setup.
Our evaluation for point cloud classification was carried out on the ModelNet40 dataset [Wu et al., 2015], which consists of 40 object categories and a total of 12,311 3D models. Following prior studies [Wang et al., 2019, Deng et al., 2021], we allocated 9,843 models for training and 2,468 models for testing in the classification task. For each model, 1,024 points were uniformly sampled from its mesh surface, using only the xyzxyz coordinates of the sampled points. We apply on-the-fly rotation augmentation during training, following Esteves et al. [2018], Deng et al. [2021], such that the dataset size remains unchanged. At test time, each example is rotated by an arbitrary rotation. Note that on-the-fly augmentation essentially changes the training data distribution during the learning process. The purpose of comparing under rotation protocols is to assess a model’s invariance to rotational changes.
Competing Baselines.
For point cloud classification tasks, we compare our anisotropic geometric method with point cloud approaches, including PointNet and DGCNN architectures, as well as equivariant models based on the vector neuron framework, i.e., VN-PointNet and VN-DGCNN. We further include canonicalization baselines, CN-PointNet and CN-DGCNN, and traditional augmentation baselines in which the training set is expanded with pre-generated rotations (PointNet-Aug, DGCNN-Aug) with a factor of five ($\times 5$). The experimental results of PointNet, DGCNN, VN-PointNet, and VN-DGCNN are taken from Wang et al. [2019], Deng et al. [2021].
Hyperparameters.
We follow the published hyperparameters and training protocol of PointNet and DGCNN. For PointNet, we uses identical channel widths to PointNet (64, 64, 64, 128, 1024). We use Adam optimizer with learning rate 0.001 and batch size 32 with a weight decay $1\times 10^{-4}$ and dropout 0.3. For DGCNN, each input comprises 1,024 uniformly sampled points, and the k-NN graph uses neighborhood size $k=20$. DGCNN uses four EdgeConv layers (with per-layer MLPs of sizes 64, 64, 128, 256). We train with stochastic gradient descent (initial learning rate 0.1) and apply a cosine annealing schedule of 0.001. Training runs for 250 epochs with a batch size of 32, and we use a dropout rate of 0.5 in the fully connected layers.
Appendix G Additional Results
In this section, we present additional results of our adaptive canonicalization, including experimental trade-offs of sampling-based and optimization-based construction, anisotropic nonlinear spectral filters for node-level representation, and out-of-sample rotation generalization for point clouds.
G.1 Sampling-based VS Optimization-based Implementation
To evaluate the trade-offs between sampling-based and our sample-and-refine (optimization-based) implementation, we conduct experiments on the TUDataset graph classification benchmarks. Tab. LABEL:tab:sampling_or_optimization reports the classification performance of TUDaset under sampling-based and optimization-based adaptive caninocalization. We see that the optimization-based implementation consistently performs better than the sampling-based one. While increasing the sampling candidates (from $1\times$ to $5\times$ or $10\times$) improves the performance, the sample-and-refine strategy is more memory-efficient than massive sampling approaches. Rather than storing and evaluating hundreds of rotation matrices simultaneously, it processes a smaller working set through iterative refinement, reducing memory pressure [Li et al., 2022]. In terms of computation time, the sampling-based method grows linearly with the number of candidates, while the optimization-based method add a small overhead to the inner steps. We see that in practice, a modest refinement (a few steps) surpasses the accuracy of large sampling budgets at a lower time, offering a better accuracy-time trade-off.
Table 5: Comparison of sampling-based vs. optimization-based adaptive canonicalization. Classification accuracy across TUDataset. Sampling methods use different numbers of random candidates, while the optimization approach combines sampling with local refinement via gradient descent.
MUTAG PTC ENZYMES PROTEINS NCI1
[…]
G.2 Node-Level Representation Anisotropic Nonlinear Filters
We introduce the adaptive canonicalization applied to spectral graph neural networks for learning graph-level representation in Sec. 3.1 and App. E.1. Importantly, the adaptive canonicalization can also be applied to node-level representation. We describe this construction below.
In this setting, we continue to consider a classification task with $D$ classes. For each class $d$, we denote its corresponding class prior $h_{d}$. The class-$d$ orientation for node representation is obtained by optimizing
$\mathbf{U}^{\square(d)}=\bigoplus_{k}{\mathbf{U}}_{k}^{\square(d)}=\underset{\mathbf{U}^{(d)}_{1},\ldots,\mathbf{U}^{(d)}_{B}}{\arg\max}\;h_{d}\left(\Psi_{d}\left(\sum_{i}\left({\mathbf{U}}^{(d)*}\phi\left(\widetilde{\mathbf{X}}^{(d)}_{1},\ldots,\widetilde{\mathbf{X}}^{(d)}_{B}\right)\right)_{k}{\mathbf{V}}_{k}^{\top}\right)\right).$
Here, the node-level representation proceeds by mapping the input signal to the spectral domain [Mallat, 2002] in a band-wise manner with an oriented basis within each band’s eigenspace, and performing a synthesis step that transforms the learned coefficients back to the node domain.
As discussed in App. D, on a square grid, each $x$ Fourier mode has a corresponding $y$ Fourier mode of the same response. Therefore, standard spectral methods are inherently isotropic as they cannot distinguish between horizontal and vertical directional information. On the other hand, adaptive canonicalization with synthesis in the node-level setting is anisotropic. Within each degenerate eigenspace, our method can learn distinct orientations, and the subsequent synthesis step maps these anisotropic spectral coefficients back to nodes. The resulting spatial operator can therefore implement any directional filter that a convolutional neural network can achieve [Shannon, 2006, LeCun and Bengio, 1998, Freeman et al., 1991, Dagès et al., 2024]. We leave a detailed exploration of such node-level tasks for future work.
However, in graph-level tasks, the canonicalized node-level embeddings can serve as the intermediate representation from which graph-level features are derived. Specifically, after applying adaptive canonicalization and synthesis at the node level, the resulting node embeddings can be aggregated through standard pooling operations to have a graph-level representation. We evaluate the node-to-graph construction on TUDataset for graph classification tasks. The results are summarized in Tab. LABEL:tab:graph_classification_by_node_representation. We see that the node-to-graph construction achieves performance closely aligned with, and in some cases approaching, that of the direct graph-level canonicalization. We attribute the slightly worse performance to the potential pooling loss.
Table 6: Graph classification performance on TUDataset using adaptive canonicalization. Comparison between direct graph-level representations (Graph) and node-to-graph constructions (Node-to-graph).
MUTAG PTC ENZYMES PROTEINS NCI1
[…]
G.3 Out-of-Sample Rotation Generalization in Point Clouds
We adopt the $z/\mathcal{SO}(3)$ protocol [Esteves et al., 2018, Deng et al., 2021]: training with on-the-fly azimuthal rotations ($z$-axis) augmentation, and evaluation applies under arbitrary rotations to each test shape. In this setting, we assess out-of-sample rotation generalization by constraining training data rotations while testing on the full rotation group. The classification performance on ModelNet40 under $z/\mathcal{SO}(3)$ protocol is reported in Tab. 7. Standard PointNet and DGCNN drop sharply under this shift. Equivariant vector-neuron variants recover much of the loss, and canonicalization baselines are comparable. Our adaptive canonicalization outperforms both equivariant architecture and canonicalization baselines in both backbones.
Table 7: Classification accuracy on ModelNet40 under $z/\mathcal{SO}(3)$ protocol.
PointNet DGCNN VN-PointNet VN-DGCNN CN-PointNet CN-DGCNN AC-PointNet AC-DGCNN
Accuracy 19.6 33.8 77.5 89.5 79.6 88.8 81.4 91.8